# MIRI spectrophotometry of GN-z11: Detection and nature of an optical red continuum component

### v1

@@
We present new MIRI F560W, F770W and F1000W imaging of the galaxy GN-z11 at a redshift of 10.603. We report a significant detection (14) in the F560W and F770W images, and a marginal detection (3.2) in the F1000W filter. The new MIRI observations cover the optical-red spectral range, and significantly extend previous NIRCam wavelength coverage from rest-frame 0.38 m up to 0.86 m. In this work, we analyse the spectral energy distribution (SED) combining this new MIRI imaging data with archival NIRSpec/Prism and MRS spectroscopy, and NIRCam imaging, i.e. covering the rest-frame 0.12 0.86 m. New constraints such as the equivalent widths of the strong optical lines ([O III] 5008, H and H) and the continuum emission at rest-frame 0.48m, 0.66m, and 0.86m, free from emission line contributions, are presented. The continuum emission shows a flat energy distribution, in , up to 0.5 m, compatible with the presence of a mixed stellar population of young (4 1 Myr) and mature (63 23 Myr) stars that also account for the [O III], H and H emission lines. The continuum at rest-frame 0.66 m shows a 36 3% flux excess above the predicted flux for a mixed stellar population, pointing to the presence of an additional source contributing at these wavelengths. This excess increases to 91 28% at rest-frame 0.86 m, although with a large uncertainty due to the marginal detection in the F1000W filter. We consider that hot dust emission in the dusty torus around a type 2 AGN could be responsible for the observed excess. Alternatively, this excess could be due to hot dust emission or to a photoluminiscence dust process (Extended Red Emission, ERE) under the extreme UV radiation field, as observed in some local metal-poor galaxies and in young compact starbursts. The presence of a type 1 AGN is not supported by the observed SED as the hot dust emission in luminous high- QSOs contributes at wavelengths above rest-frame 1 m, and an additional ad-hoc red source would be required to explain the observed flux excess at 0.66 and 0.86 m. Additional deep MIRI imaging covering the rest-frame near-IR are needed to confirm the flux detection at 10 m with higher significance, and to discriminate between the different hot dust emission in the extreme starburst and AGN scenarios with MIRI imaging at longer wavelengths.
@@
GN-z11 was firstly identified as a 10 Lyman-break galaxy candidate using multi-colour imaging from HST CANDELS (Bouwens+10). A subsequent HST grism spectroscopic analysis revealed a flux dropout identified as the Lyman break by Oesch+16, yielding a redshift value of 11.09. The arrival of JWST has enabled the study of this object at optical and near-IR wavelength with unprecedented sensitivity and spatial resolution. In particular, the JWST Advanced Deep Extragalactic Survey (JADES; Eisenstein+23) derived a redshift = 10.6034 0.0013 using multiple emission lines based on NIRSpec/Prism data (Bunker+23). Besides, this NIRSpec spectrum has revealed the presence of a high ionization field (log() 2.3) and a super-solar abundance of nitrogen ([N/O] 0.49).Recent JWST NIRCam imaging have shown that GN-z11 UV-emission can be depicted as the combination of a point-source and an extended component, showing an extremely compact morphology (combined of 64 20 pc) and large UV luminosity ( = 21.58 0.02; Tacchella+23_GNz11).
@@
JWST multi-wavelength photometry and spectroscopy result on star-formation rates (SFRs) and stellar masses between 20 30 yr-1 and log(/ ) = 8.79.1 (Bunker+23; Tacchella+23_GNz11), respectively. The relatively large stellar mass at only 420 Myr after the Big Bang suggests a rapid build-up of stellar mass, which has been also found in other JWST spectroscopically confirmed galaxies at 10 (Robertson+23; Curtis-Lake+23; Carniani+24; Naidu+25). The exceptionally UV brightness of this object allows its observation using ground-based facilities. Recent NOEMA observations (Fudamoto2024) have determined upper limits on the [C II]158 m emission and 160 m dust continuum yielding log(/ ) 9.3 and log(/ ) 6.9. These results are consistent with a negligible dust attenuation and the blue colour ( = -2.4) obtained from a SED-fitting analysis (Tacchella+23_GNz11).
@@
Using medium resolution (R = 1000) NIRSpec observations, Maiolino+24_BH present several pieces of evidences favouring the scenario where GN-z11 is a Narrow Line Seyfert 1 (NLS1) with a black hole mass of log() = 6.2 0.3 accreting material at a large Eddington ratio ( = 5.5). The authors found the presence of high-excitation ( 60 eV) emission lines in the UV spectrum (e.g., [Ne IV]2422,2244) which are usually present in AGN (LeFevre+19; Terao+22). In addition, the high critical density ( 105 cm-3) and broad (FWHM 1200 km s-1) emission lines revealed by the NIRSpec data are typically observed in the broad-line regions (BLRs) of AGNs. The presence of an outflow powered by this AGN was also inferred from a high equivalent width and blue-shifted absorption of a C IV doublet (Maiolino+24_BH). The small and compact size of the BLR linked to this potential AGN would justify the large [N/O] of GN-z11 through a very rapid chemical enrichment driven by only a few SNe. In addition, a blue bump has been observed in the UV using NIRSpec/Prism data (Ji+25). This bump has been proposed to be linked to dense Fe II-emitting clouds ( 1011 cm-3) infalling at 3000 km s-1. GN-z11 hosting an AGN and the presence of tentative nearby companions detected as Ly blobs make this region a possible candidate for being the core of a protocluster with a halo mass log(/ ) = 10.4 (Scholtz+24).
@@
However, alternative scenarios have been proposed to justify the overabundance of nitrogen in GN-z11, such as the presence of Wolf-Rayet (WR) stars (Senchyna2024) or super-massive stars in a proto-globular cluster (Charbonnel2023). In addition, Bhatt2024 show that cosmological simulations do not support a super-Eddington type 1 AGN scenario for GN-z11. The analysis of the H and [O III] 5008 lines observed with MIRI/MRS carried out by Alvarez-marquez+25 does not support the presence of an accreting black hole dominating the optical emission lines and continuum in GN-z11. In the case H and the optical continuum are entirely produced by the BLR, the accretion rate and bolometric luminosity would be unrealistically large, but the black hole mass would be smaller (log(/ ) = 5.77) than the one inferred by UV lines in Maiolino+24_BH. In addition, and / were found inconsistent with the low- AGN relations (Ho2001; Greene+05b).
@@
To date, the majority of studies on GN-z11 have relied on UV data. However, we still lack important information at optical wavelengths, where the contribution from the young stellar populations begin to decrease and those from more evolved populations, or redder emission sources, is expected to become more prominent. So far, only Alvarez-marquez+25 has covered the optical wavelengths, using MIRI/MRS data, detecting [O III] 5008 and H emission consistent with star formation. This paper presents the first mid-infrared F560W and F770W images of GN-z11, tracing its optical emission at rest-frame 0.5 and 0.66 m, respectively. We combine these images with ancillary spectro-photometric data covering from the rest- 0.15 up to 0.86 m to trace the UV to optical emission continuum of this object and extend the analysis of its SED up to optical-red wavelengths. The paper is organised as follows. Section 2 presents the new JWST/MIRI images used in this work and the ancillary JWST data. Section 3 describes the continuum-only extraction and SED fitting analysis carried out. The main results drawn from the analysis and their discussion are presented in Section 4. Finally, Section 5 summarises the main conclusions and results of this work. Throughout this paper we assume a Chabrier initial mass function (IMF, Chabrier+03) and a flat CDM cosmology, with = 0.31 and H0 = 67.7 km s-1 Mpc-1 (PlanckCollaboration18VI). For this cosmology, 1 arcsec corresponds to 4.08 kpc at = 10.603 while the luminosity distance is = 113.4 Gpc.
@@ 2.1 JWST/MIRI data and calibration
The JWST images of the GN-z11 were obtained on December 8, 2023 using the MIRI imager (MIRIM, Bouchet+15) with the F560W and F770W filters, as part of the cycle 1 JWST program ID 1264 (PI: Colina, L.). The observations were performed with the FASTR1 read-out mode in a five- and seven-point dither, for the F560W and F770W respectively, with a medium-size cycling pattern. The total on-source integration time corresponds to 2012 and 5866 seconds distributed in 10 and 28 integrations, respectively for F560W and F770W.
These MIRIM images have been calibrated using the JWST pipeline (v1.12.0) with the context 1170 of the Calibration Reference Data System (CRDS). This CRDS includes photometric calibrations considering the temporal evolution and aperture corrections taking into account the point-spread function (PSF) cruciform. In addition to the general procedure, further steps have been applied to correct for striping and background gradients (see more details in Alvarez-Marquez+23; Perez-Gonzalez+24; Ostlin+25). The process concluded with the creation of final MIRI images dithered to a pixel scale of 0.06/pixel. According to the JWST documentation111https://jwst-docs.stsci.edu/jwst-mid-infrared-instrument/miri-performance/miri-point-spread-functions, the spatial resolutions for these images are FWHM = 0.207 and 0.269, for the F560W and F770W, respectively. Following this calibration, we obtain a 3 depth of 26.67 and 26.96 mag for F560W and F770W for unresolved point sources, respectively.
@@ 2.2 Ancillary JWST data: photometry and spectroscopy
In addition to the MIRI F560W and F770W images, this work takes advantage of the large amount of high-quality JWST data available for GN-z11 covering a large wavelength range. To trace the near-IR counterpart of this galaxy, we use the F1000W and F2100W images available from the MEOW survey (PID 5407, PI: G. Leung) which covers up to rest-frame 2 m. These observations were performed with a total exposure time of 722 and 3086 s reaching a 3 level of 230 nJy and 2.6 Jy (i.e., 25.5 and 22.9 mag, respectively). In addition, the 11 NIRCam images available from JADES DR3 (Eisenstein+23; Bunker+24; DEugenio+25), with median 3 depths of 5 nJy (i.e., 29.6 mag, DEugenio+25), are used to trace the UV continuum from rest-frame 0.1 up to 0.4 m. We performed aperture photometry as explained in Sect. 3.1 for all the NIRCam and MIRI images available.
We have also incorporated JWST ancillary spectroscopic data for this object. In particular, we consider the fluxes from the UV lines detected on the NIRSpec/Prism data presented in Bunker+23 after a re-normalization to match the NIRCam photometry (see Sect. 3.1 and App. A). In addition to the UV lines from NIRSpec, we also use the flux from the optical lines present in the MRS data and extracted by Alvarez-marquez+25. The MRS spectra contain the H + [O III] and H lines which fall into the MIRI/F560W and F770W wavelength ranges, respectively. We do not apply any normalization as the MRS line fluxes were corrected for aperture losses using the MRS PSFs (Argyriou+23).
@@ 3 Analysis
3.1 Aperture photometry and NIRSpec renormalization
@@
We extract the GN-z11 photometry from the available NIRCam and MIRI images considering circular apertures, centred on the F560W emission peak. The size of these apertures (i.e., = 0.5) has been chosen to fully encompass both the UV and the optical emission. Before extracting the fluxes from NIRCam, we matched the PSF of these images to the spatial resolution of MIRI/F560W. At MIRI resolution, we consider GN-z11 as a point-like source and therefore we applied the aperture corrections derived from the empirical PSFs presented in Libralato+24. At = 0.5, these corrections are 1.29, 1.33, 1.54 for MIRI/F560W and the NIRCam images, MIRI/F770W and MIRI/F2100W, respectively. To optimise the S/N on the F1000W detection we have assumed an aperture of = 0.3, applying an aperture correction factor of 1.71. The associated uncertainty of each photometric measurement is defined as the standard deviation of the local background around GN-z11 (1″¡ ¡ 3″), masking all the additional objects present in the field-of-view (FoV). In this work, we have re-scaled the MIRI background noise by a factor of 2.24 to take into account the correlated noise induced when drizzling the individual observations (see Ostlin+25, for details). We adopt a 3 upper limit for the F2100W image, which does not present significant detection. Table 1 presents the flux values measured from the NIRCam and MIRI images. The S/N is generally high (10) for all photometric values, with the exception of F1000W, in which GN-z11 is detected at a 3 level.
We also use the NIRCam fluxes to renormalise the NIRSpec spectra, correcting from potential slit loses. In brief, we matched the NIRSpec/Prism spectra to the NIRCam photometry multiplying by a factor of 1.325. A detailed description of the process is presented in Appendix A.
@@ 3.2 Continuum-only fluxes and EWs
Since the optical emission lines detected in the MRS (i.e., [O III] 5008, H) fall within the F560W and F770W filters, respectively, their contribution can be subtracted from the broadband fluxes, allowing us to estimate the underlying continuum at 5.6 and 7.7 m, and derive the equivalent width (EW) of these emission lines.
In this work, we have used the [O III] 5008 and H fluxes presented by Alvarez-marquez+25 and the transmission curves for the F560W and F770W filters333Transmission curves are obtained from the SVO Filter Profile Service. For accuracy, we have included the [O III] 4960 and H contributions using [O III] 5008/[O III] 4960 = 3 and H/H = 2.86, considering case B and with negligible extinction (Bunker+23; Alvarez-marquez+25). Assuming log() = 3 and 0.2 , the potential contributions from [N II6549,6585 and [S II]6718,6732 to the F770W flux are minimal (5, Curti+20).
In addition, we have extended the derivation of the continuum levels to UV wavelengths by using the NIRCam fluxes and the NIRSpec/Prism re-normalised emission lines presented in Bunker+23. Finally, we estimate the continuum in MIRI/F1000W based on the [S III] 9071,9533/H = 0.17 ratio derived from Kewley+02, assuming 0.2 and log() = 3. Although F1000W is slightly narrower than F770W, the [S III] 9071,9533 flux ( 12 nJy) would contribute only up to 5 of the total F1000W photometry. The computed continuum-only fluxes for the NIRCam and MIRI filters are listed in Table 1.
Alvarez-marquez+25 estimate an upper limit of a possible broad component in H of 2 3 10-18 erg s-1 cm-2. If we consider this broad component, the F560W and F770W continuum levels would decrease by 6 9 nJy and 18 27 nJy, respectively, assuming a ratio H/H = 2.86 for the broad component. Based on the MIRI/F560W and F770W continuum-only fluxes, we derive rest-frame EW([O III] 5008) = 786 81 and EW(H) = 547 72, considering only the narrow line contribution on H.

### v2

@@
We present new MIRI F560W, F770W, and F1000W imaging of the galaxy GN-z11 at a redshift of 10.603. We report a significant detection (14) in the F560W and F770W images, and a marginal detection (3.2) in the F1000W filter. The new MIRI observations cover the optical-red spectral range and significantly extend previous NIRCam wavelength coverage from rest-frame 0.38 m up to 0.86 m. In this work, we analyse the spectral energy distribution (SED) combining this new MIRI imaging data with archival NIRSpec/Prism and MRS spectroscopy, and NIRCam imaging, i.e. covering the rest-frame 0.12 0.86 m. New constraints such as the equivalent widths of the strong optical lines ([O III] 5008, H and H) and the continuum emission at rest-frame 0.48m, 0.66m, and 0.86m, free of emission line contributions, are presented. The continuum emission shows a flat energy distribution, in , up to 0.5 m, compatible with the presence of a mixed stellar population of young (4 1 Myr) and mature (63 23 Myr) stars that also account for the [O III], H, and H emission lines. The continuum at rest-frame 0.66 m shows a 36 3% flux excess above the predicted flux for a mixed stellar population, pointing to the presence of an additional source contributing at these wavelengths. This excess increases to 91 28% at rest-frame 0.86 m, although with a large uncertainty due to the marginal detection in the F1000W filter. We consider that hot dust emission in the dusty torus around a type 2 active galactic nucleus (AGN) could be responsible for the observed excess. Alternatively, this excess could be due to hot dust emission or a photoluminiscence dust process (Extended Red Emission, ERE) under the extreme UV radiation field, as is observed in local metal-poor galaxies and in young compact starbursts. The presence of a type 1 AGN is not supported by the observed SED as the hot dust emission in luminous high- quasi-stellar objects (QSOs) contributes at wavelengths above rest-frame 1 m, and an additional ad hoc red source would be required to explain the observed flux excess at 0.66 and 0.86 m. Additional deep MIRI imaging covering the rest-frame near-IR is needed to confirm the flux detection at 10 m, and to discriminate between the different hot dust emission in the extreme starburst and AGN scenarios.
@@
GN-z11 was firstly identified as a 10 Lyman-break galaxy candidate using multi-colour imaging from HST CANDELS (Bouwens et al., 2010). A subsequent HST grism spectroscopic analysis revealed a flux dropout identified as the Lyman break by Oesch et al. (2016), yielding a redshift value of 11.09. The arrival of JWST has enabled the study of this object at optical and near-IR wavelength with unprecedented sensitivity and spatial resolution. In particular, the JWST Advanced Deep Extragalactic Survey (JADES; Eisenstein et al. 2023) derived a redshift of = 10.6034 0.0013 using multiple emission lines based on NIRSpec/Prism data (Bunker et al., 2023). Besides, this NIRSpec spectrum has revealed the presence of a high ionisation field (log() 2.3) and a super-solar abundance of nitrogen ([N/O] 0.49). Recent JWST NIRCam imaging has shown that GN-z11 UV emission can be depicted as the combination of a point-source and an extended component, showing an extremely compact morphology (combined of 64 20 pc) and large UV luminosity ( = 21.58 0.02; Tacchella et al. 2023a).
@@
JWST multi-wavelength photometry and spectroscopy result in star formation rates (SFRs) and stellar masses of between 20 30 yr-1 and log(/) = 8.79.1 (Bunker et al., 2023; Tacchella et al., 2023a), respectively. The relatively high stellar mass at only 420 Myr after the Big Bang suggests a rapid build-up of stellar mass, which has been also found in other JWST spectroscopically confirmed galaxies at 10 (Robertson et al., 2023; Curtis-Lake et al., 2023; Carniani et al., 2024; Naidu et al., 2025). The exceptionally UV brightness of this object allows its observation using ground-based facilities. Recent NOEMA observations (Fudamoto et al., 2024) have determined upper limits on the [C II]158 m emission and 160 m dust continuum yielding log(/) 9.3 and log(/) 6.9. These results are consistent with a negligible dust attenuation and the blue colour ( = -2.4) obtained from a spectral energy distribution (SED) fitting analysis (Tacchella et al., 2023a).
@@
Using medium-resolution (R = 1000) NIRSpec observations, Maiolino et al. (2024) present several pieces of evidence favouring the scenario in which GN-z11 is a narrow-line Seyfert 1 (NLS1) with a black hole mass of log() = 6.2 0.3 accreting material at a large Eddington ratio ( = 5.5). The authors found the presence of high-excitation ( 60 eV) emission lines in the UV spectrum (e.g. [Ne IV]2422, 2244) that are usually present in active galactic nuclei (AGNs) (Le Fèvre et al., 2019; Terao et al., 2022). In addition, the high critical density ( 105 cm-3) and broad (FWHM 1200 km s-1) emission lines revealed by the NIRSpec data are typically observed in the broad-line regions (BLRs) of AGNs. The presence of an outflow powered by this AGN was also inferred from a high equivalent width (EW) and blueshifted absorption of a C IV doublet (Maiolino et al., 2024). The small and compact size of the BLR linked to this potential AGN would justify the large [N/O] of GN-z11 through a very rapid chemical enrichment driven by only a few supernovae (SNe). In addition, a blue bump has been observed in the UV using NIRSpec/Prism data (Ji et al., 2025). This bump has been proposed to be linked to dense Fe II-emitting clouds ( 1011 cm-3) infalling at 3000 km s-1. GN-z11 hosting an AGN and the presence of tentative nearby companions detected as Ly blobs make this region a possible candidate for being the core of a protocluster with a halo mass of log(/) = 10.4 (Scholtz et al., 2024).
@@
However, alternative scenarios have been proposed to justify the overabundance of nitrogen in GN-z11, such as the presence of Wolf-Rayet (WR) stars (Senchyna et al., 2024) or super-massive stars in a proto-globular cluster (Charbonnel et al., 2023). In addition, Bhatt et al. (2024) show that cosmological simulations do not support a super-Eddington type 1 AGN scenario for GN-z11. The analysis of the H and [O III] 5008 lines observed with MIRI/MRS carried out by Álvarez-Márquez et al. (2025) does not support the presence of an accreting black hole dominating the optical emission lines and continuum in GN-z11. In the case in which H and the optical continuum are entirely produced by the BLR, the accretion rate and bolometric luminosity would be unrealistically large, but the black hole mass would be smaller (log(/) = 5.77) than the one inferred by UV lines in Maiolino et al. (2024). In addition, and / were found to be inconsistent with the low- AGN relations (Ho et al., 2001b; Greene and Ho, 2005).
@@
To date, the majority of studies on GN-z11 have relied on UV data. However, we still lack important information at optical wavelengths, where the contributions from the young stellar populations begin to decrease and those from more evolved populations, or redder emission sources, are expected to become more prominent. So far, only Álvarez-Márquez et al. (2025) have covered the optical wavelengths, using MIRI/MRS data, detecting [O III] 5008 and H emission consistent with star formation. This paper presents the first mid-infrared F560W and F770W images of GN-z11, tracing its optical emission at rest-frame 0.5 and 0.66 m, respectively. We combine these images with ancillary spectro-photometric data covering the rest-frame 0.15 up to 0.86 m to trace the UV to optical emission continuum of this object and extend the analysis of its SED up to optical-red wavelengths. The paper is organised as follows. Section 2 presents the new JWST/MIRI images used in this work and the ancillary JWST data. Section 3 describes the continuum-only extraction and SED fitting analysis carried out. The main results drawn from the analysis and their discussion are presented in Section 4. Finally, Section 5 summarises the main conclusions and results of this work. Throughout this paper we assume a Chabrier initial mass function (IMF, Chabrier 2003) and a flat cold dark matter (CDM) cosmology, with = 0.31 and H0 = 67.7 km s-1 Mpc-1 (Planck Collaboration et al., 2020). For this cosmology, 1 arcsec corresponds to 4.08 kpc at = 10.603, while the luminosity distance is = 113.4 Gpc.
@@ 2.1 JWST/MIRI data and calibration
The JWST images of the GN-z11 were obtained on December 8, 2023 using the MIRI imager (MIRIM, Bouchet et al., 2015) with the F560W and F770W filters, as part of the cycle 1 JWST program ID 1264 (PI: Colina, L.). The observations were performed with the FASTR1 read-out mode in a five- and seven-point dither, for F560W and F770W, respectively, with a medium-size cycling pattern. The total on-source integration time corresponds to 2012 and 5866 seconds distributed in 10 and 28 integrations, respectively, for F560W and F770W.
These MIRIM images have been calibrated using the JWST pipeline (v1.12.0) with the context 1170 of the Calibration Reference Data System (CRDS). This CRDS includes photometric calibrations considering the temporal evolution and aperture corrections taking into account the point spread function (PSF) cruciform. In addition to the general procedure, further steps have been applied to correct for striping and background gradients (see more details in Álvarez-Márquez et al. 2023; Pérez-González et al. 2024a; Östlin et al. 2025). The process concluded with the creation of final MIRI images dithered to a pixel scale of 0.06/pixel. According to the JWST documentation111https://jwst-docs.stsci.edu/jwst-mid-infrared-instrument/miri-performance/miri-point-spread-functions, the spatial resolutions for these images are FWHM = 0.207 and 0.269, for the F560W and F770W, respectively. Following this calibration, we obtained a 3 depth of 26.67 and 26.96 mag for F560W and F770W for unresolved point sources, respectively.
@@ 2.2 Ancillary JWST data: Photometry and spectroscopy
In addition to the MIRI F560W and F770W images, this work takes advantage of the large amount of high-quality JWST data available for GN-z11 covering a large wavelength range. To trace the near-IR counterpart of this galaxy, we used the F1000W and F2100W images available from the MEOW survey (PID 5407, PI: G. Leung) which covers up to rest-frame 2 m. These observations were performed with a total exposure time of 722 and 3086 s reaching a 3 level of 230 nJy and 2.6 Jy (i.e. 25.5 and 22.9 mag, respectively). In addition, the 11 NIRCam images available from JADES DR3 (Eisenstein et al., 2023; Bunker et al., 2024; D’Eugenio et al., 2025), with median 3 depths of 5 nJy (i.e. 29.6 mag, D’Eugenio et al. 2025), were used to trace the UV continuum from rest-frame 0.1 up to 0.4 m. We performed aperture photometry as is explained in Sect. 3.1 for all the NIRCam and MIRI images available.
We also incorporated JWST ancillary spectroscopic data for this object. In particular, we considered the fluxes from the UV lines detected on the NIRSpec/Prism data presented in Bunker et al. (2023) after a re-normalisation to match the NIRCam photometry (see Sect. 3.1 and App. A). In addition to the UV lines from NIRSpec, we also used the flux from the optical lines present in the MRS data and extracted by Álvarez-Márquez et al. (2025). The MRS spectra contain the H + [O III] and H lines that fall into the MIRI/F560W and F770W wavelength ranges, respectively. We did not apply any normalisation as the MRS line fluxes were corrected for aperture losses using the MRS PSFs (Argyriou et al., 2023).
@@ 3 Analysis
3.1 Aperture photometry and NIRSpec re-normalisation
@@
We extracted the GN-z11 photometry from the available NIRCam and MIRI images considering circular apertures, centred on the F560W emission peak. The size of these apertures (i.e. = 0.5) has been chosen to fully encompass both the UV and the optical emission. Before extracting the fluxes from NIRCam, we matched the PSF of these images to the spatial resolution of MIRI/F560W. At MIRI resolution, we consider GN-z11 to be a point-like source, and therefore we applied the aperture corrections derived from the empirical PSFs presented in Libralato et al. (2024). At = 0.5, these corrections are 1.29, 1.33, and 1.54 for MIRI/F560W and the NIRCam images, MIRI/F770W and MIRI/F2100W, respectively. To optimise the signal-to-noise (S/N) on the F1000W detection we assumed an aperture of = 0.3, applying an aperture correction factor of 1.71. The associated uncertainty of each photometric measurement was defined as the standard deviation of the local background around GN-z11 (1″¡ ¡ 3″), masking all the additional objects present in the field of view (FoV). In this work, we have re-scaled the MIRI background noise by a factor of 2.24 to take into account the correlated noise induced when drizzling the individual observations (see Östlin et al. 2025, for details). We adopted a 3 upper limit for the F2100W image, which does not present significant detection. Table 1 presents the flux values measured from the NIRCam and MIRI images. The S/N is generally high (10) for all photometric values, with the exception of F1000W, in which GN-z11 is detected at a 3 level (see Fig. 1).
We also used the NIRCam fluxes to re-normalise the NIRSpec spectra, correcting from potential slit loses. In brief, we matched the NIRSpec/Prism spectra to the NIRCam photometry, multiplying by a factor of 1.325. A detailed description of the process is presented in Appendix A.
@@ 3.2 Continuum-only fluxes and EWs
Since the optical emission lines detected in the MRS (i.e. [O III] 5008, H) fall within the F560W and F770W filters, respectively, their contribution can be subtracted from the broadband fluxes, allowing us to estimate the underlying continuum at 5.6 and 7.7 m and derive the EW of these emission lines.
In this work, we have used the [O III] 5008 and H fluxes presented by Álvarez-Márquez et al. (2025) and the transmission curves for the F560W and F770W filters333Transmission curves are obtained from the SVO Filter Profile Service. For accuracy, we have included the [O III] 4960 and H contributions using [O III] 5008/[O III] 4960 = 3 and H/H = 2.86, considering case B and with negligible extinction (Bunker et al., 2023; Álvarez-Márquez et al., 2025). Assuming log() = 3 and 0.2 , the potential contributions from [N II]6549, 6585 and [S II]6718, 6732 to the F770W flux are minimal (5, Curti et al. 2020).
In addition, we have extended the derivation of the continuum levels to UV wavelengths by using the NIRCam fluxes and the NIRSpec/Prism re-normalised emission lines presented in Bunker et al. (2023). Finally, we estimated the continuum in MIRI/F1000W based on the [S III] 9071,9533/H = 0.17 ratio derived from Kewley and Dopita (2002), assuming 0.2 and log() = 3. Although F1000W is slightly narrower than F770W, the [S III] 9071,9533 flux ( 12 nJy) would contribute only up to 5 of the total F1000W photometry. The computed continuum-only fluxes for the NIRCam and MIRI filters are listed in Table 1.
Álvarez-Márquez et al. (2025) estimate an upper limit of a possible broad component in H of 2 3 10-18 erg s-1 cm-2. If we consider this broad component, the F560W and F770W continuum levels would decrease by 6 9 nJy and 18 27 nJy, respectively, assuming a ratio H/H = 2.86 for the broad component. Based on the MIRI/F560W and F770W continuum-only fluxes, we derived rest-frame EW([O III] 5008) = 786 81 and EW(H) = 547 72, considering only the narrow line contribution to H.